OpenStudy (anonymous):

simplify:

Parth (parthkohli):

Nothing can't be simplified. Check it again.

OpenStudy (anonymous):

\[5\sqrt{3} (2\sqrt{3} + 6\sqrt{19}\]

OpenStudy (cwtan):

hahaahaha @ParthKohli

OpenStudy (anonymous):

it doesnt let me use equation when i write questions :l only when i comment

OpenStudy (cwtan):

if there is a like button here i will click it

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow (5\sqrt{3} \times 2\sqrt3) + (5\sqrt{3} \times 6\sqrt{19}) }\)

Parth (parthkohli):

Distributive property it is.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 30 + (30\sqrt{57}) }\)

OpenStudy (anonymous):

@Brad1996, aren't you starting to get an intuition for solving radicals now?

Parth (parthkohli):

Yeah, no more question please lol

Parth (parthkohli):

questions*

OpenStudy (anonymous):

im working on like 4 paper rite now :l

Parth (parthkohli):

Omg

OpenStudy (anonymous):

trying to answer one and ask questions for the other on OS

OpenStudy (cwtan):

4 papers can be easily done LOL

OpenStudy (anonymous):

yea but i wanna be done in like an hr not 4 hrs :d

Parth (parthkohli):

Just 4 papers? Y U NO GIVE 10 MORE?

OpenStudy (cwtan):

Is there a like button for the comment above? I want to click it

OpenStudy (anonymous):

what do u mean?

OpenStudy (cwtan):

Like the comment ~ haha

OpenStudy (anonymous):

dont think so

OpenStudy (anonymous):

woohoo question 47 out of 100 XD

OpenStudy (cwtan):

400 /60minute approx 6 question per minute

OpenStudy (anonymous):

dude it takes me like 5minutes for 1 question lol

OpenStudy (cwtan):

wow u need to answer 1 question in 10 second~~~~ that was rocket

OpenStudy (anonymous):

48 out of 100 :d

OpenStudy (anonymous):

can u answer this question please? Last 1 on my 2nd paper.. :d

OpenStudy (cwtan):

God Bless You